# Reinforcement Learning-based Adaptive Path Selection for Programmable Networks

### v1

@@ V Experimental Evaluation
Figure 2: Proof-of-Concept implementation of the SLA-based path selection system
V-A Experimental Results
This section evaluates the RL agent’s response to real-time congestion in the PoC topology. We assess how effectively the agent adapts its path selection based on queue length and dequeueing delay, demonstrating the SLA-based update mechanism’s ability to guide stable and efficient forwarding decisions.
@@ V-A Experimental Results
Figure 3: Evolution of queue metrics and path selection over time for learning rate $\alpha=0.5$. The RL agent adapts to congestion by shifting traffic to the less loaded path.
@@
Figure 4: Performance evaluation: (a) Learning rate impact on convergence time, (b) average throughput comparison between baseline forwarding and SLA-based forwarding with $\alpha=0.5$

### v2

@@ V-A Setup and Implementation
Figure 2: Proof-of-Concept topology with five BMv2 switches and two hosts, including decision node ($S_{1}$), sink ($S_{3}$), and collector ($S_{5}$).
V-B Results
We now present the experimental results of our PoC, focusing on how the RL agent responds to real-time congestion. The evaluation examines the agent’s ability to adapt its forwarding decisions based on queue length and dequeueing delay, highlighting how the SLA-based update mechanism enables both stability and responsiveness in traffic steering.
@@ V-B Results
Figure 3: Evolution of queue length, dequeueing delay, and path selection for $\alpha=0.5$.
@@
Figure 4: Performance evaluation: (a) Impact of learning rate on convergence time, (b) average throughput comparison between baseline forwarding and SLA-based forwarding with $\alpha=0.5$